• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name saw

    Log entry time 10:29:52 on March 14, 2005

    Entry number 141463

    keyword=HV EPICS control

    The BB/ND HV EPICS control was fixed this morning. The problem turned out to be an NFS configuration problem on accelerators computers.

    I also reterminated both ends of the ethernet cable that goes to the IOC, using a different set of pairs in the wire. (The cable has 4 pairs, ethernet needs 2 pairs.) Turns out this was not necessary, but didn't hurt anything.